How many years will it take $100 to triple to $3.00 at an annual rate of 596 & Select one O a 28.5 O 0.615 O c.225 TO
Answer #1

Answer:

Given ,

Future Value = $ 3.00

Present Value = $ 1.00

Annual Rate = 5%

Future Value = Present Value * (1 + Rate)^ n

3 = 1* (1 + 0.05)^n

log 3 = n log (1.05)

n =\frac{log3}{log (1.05)} = \frac{0.4771 }{ 0.0211}

n = 22.611

Hence, It will take approximately 22 years 6 months for the money to triple.